A firm is considering a project that will generate perpetual after-tax cash flows of $17,000 per year beginning next year. The project has the same risk as the firm’s overall operations and must be financed externally. Equity flotation costs 15 percent and debt issues cost 4 percent on an after-tax basis. The firm’s D/E ratio is 0.5.


What is the most the firm can pay for the project and still earn its required return? (Do not round intermediate calculations. Round your final answer to the nearest whole dollar.)
Answer #1

wD = D / [D + E] = 0.5 / [1 + 0.5] = 0.3333

wE = E / [D + E] = 1 / [1 + 0.5] = 0.6667

WACC =[wD * kD] + [wE * kE]

= [0.3333 * 4%] + [0.6667 * 15%] = 1.33% + 10% = 11.33%

PV = Perpetual after-tax CFs / WACC = $17,000 / 0.1133 = $150,000